# Whitney Approximation: domains and bounds

### v2

@@ Introduction
In this paper we let $a$, $b$, $s$, $t$ range over $\mathbb{R}$}, $m$, $n$ over ${\mathbb{N}}=\{0,1,2,\dots\}$, $r$ over $\mathbb{N}$$\cup\{\infty\}$}, and $z$ over $\mathbb{C}$}. Let $I=(\alpha,\beta)$, where $\infty\leqslant\alpha<\beta\leqslant\infty$, be an open interval. Every $\mathcal{C}^{r}$-function $I\to{\mathbb{R}}$} and its derivatives can be arbitrarily well approximated by an analytic function and its derivatives; more precisely:
@@
This theorem is implicit in the proof of a (more general) lemma in Whitney’s seminal 1934 paper on extending multivariate differentiable functions [25]. In [3, Appendix A] we included a self-contained proof of this fact, which makes it apparent that there is an open subset $U=U_{I}$ of $\mathbb{C}$} which contains $I$, independent of $f$, such that the function $g$ in the conclusion of the theorem can be taken as the restriction of a holomorphic function $U\to{\mathbb{C}}$}. If $I$ is bounded, then the open set $U$ obtained through the proof of Whitney’s theorem as in loc. cit. is also bounded. (See [3, Corollary A.4].) However, if $I={\mathbb{R}}$, then this procedure yields $U={\mathbb{C}}$, so $g$ can be taken as the restriction of an entire function. Independently of Whitney, this fact had also been discovered by Hoischen [14], and for $r=0$ had already been shown by Carleman [8] in 1927 and for $r=1$ by Kaplan [15] in 1955. (See [20] for the history of such approximation theorems, going back to Weierstrass.) By a theorem of Gauthier and Kienzle [13, Theorem 1.3], in the case $r<\infty$} one can always take $U=I\cup({\mathbb{C}}\setminus{\mathbb{R}})$. Our first main result strengthens this fact by showing that it is possible to choose $U={\mathbb{C}}\setminus\{\alpha,\beta\}$ (even when $r=\infty$):
@@
If $\lambda_{m}(g)<\infty$ for some $m$, then $g$ is said to have finite index, and in this case the smallest such $m$ is called the index of $g$. Note that if $g$ has finite index $m$, then $m\geqslant 1$}. By convention, constant functions $\mathbb{C}$$\to{\mathbb{C}}$ have index 0. The entire functions of index $1$ are exactly the non-constant polynomial functions. The entire functions of index $\leqslant 2$ are also known as the entire functions of finite order, and  $\lambda_{2}(g)$ is called the order of $g$. (See [6, Chapter 2].) By [22, Theorem 1], if $g$ has index $m\geqslant 2$, then $\lambda=\lambda_{m}(g)$ may be computed from the Taylor coefficients $g_{n}:=g^{(n)}(0)/n!$ of $g$ at 0 as follows:
@@
is the basis of a topology on $\mathcal{S}({\mathbb{R}})$, which makes the $\mathbb{R}$}-vector space $\mathcal{S}({\mathbb{R}})$ into a topological $\mathbb{R}$}-vector space (in fact, a Fréchet space), known as the Schwartz space on $\mathbb{R}$}, which plays an important role in the theory of distributions (cf. [23]). It is well-known that the set of rapidly decreasing functions which extend to entire functions is dense in $\mathcal{S}({\mathbb{R}})$; see [23, Theorem 15.5, p. 160]. From Theorem 2 we obtain an effective version of this fact:
@@ 2. Proof of Theorem 1
Next we prove Corollary 1. Let $I$, $f$, $\varepsilon$, $\rho$ be as in the corollary. The case where $I$ is open is covered by Theorem 1, and the case where $I$ is closed and bounded is a consequence of the Weierstrass Approximation Theorem [18, p. 33]. Suppose $I=[\alpha,\beta)$} where $\infty<\alpha<\beta\leqslant\infty$. (The remaining case where $I=(\alpha,\beta]$ with $\infty\leqslant\alpha<\beta<\infty$ is treated in a similar way.) Then $\operatorname{fr}(I)=\{\beta\}$. We set $I_{1}:=(-\infty,\beta)$ and now use a well-known fact:
@@ 5. Whitney Approximation with Bounds
$\leqslant\big{(}2c(\rho(s)+1)\big{)}^{d(\rho(s)+1)(s/\sqrt{2}-1)}\cdot\lVert f\rVert_{s;\,\rho(s)+1}$

### v3

@@ Introduction
The Weierstrass Approximation Theorem regarding the approximation of a continuous function $f$ on a compact interval by polynomials was generalized to the complex domain by Mergelyan [18] in 1951, who replaced the compact interval by a compact subset $K$ of the complex plane having connected complement with the additional (necessary) condition on $f$ that it be not only continuous on $K$ but also holomorphic on the interior of $K$. (See [12, Chapter III, §2] or [23, Chapter 20] for a proof.) This powerful theorem sparked the development of the subject of complex approximation theory in the second half of the 20th century, and has had numerous applications in many branches of complex analysis. Carleman [8] had previously (1927) found another analogue of the Weierstrass theorem, in which he replaced the compact interval by the real line and polynomials by entire functions. (Actually, Carleman’s theorem can be deduced without much effort from Mergelyan’s, cf. [21, proof of Theorem 8].) Staying with the approximation of functions on the real line, the present paper is devoted to a generalization of Carleman’s theorem due to Whitney (1934).
To explain Whitney’s result, we fix some conventions which will remain in force throughout the rest of this paper: we let $a$, $b$, $s$, $t$ range over $\mathbb{R}$}, $m$, $n$ over ${\mathbb{N}}=\{0,1,2,\dots\}$, $r$ over $\mathbb{N}$$\cup\{\infty\}$}, and $z$ over $\mathbb{C}$}. Let $I=(\alpha,\beta)$, where $\infty\leqslant\alpha<\beta\leqslant\infty$, be an open interval. Each $\mathcal{C}^{r}$-function $I\to{\mathbb{R}}$} and its derivatives can be arbitrarily well approximated by an analytic function and its derivatives; more precisely:
@@
This theorem is implicit in the proof of a (more general) lemma in Whitney’s seminal 1934 paper on extending multivariate differentiable functions [27]. In [3, Appendix A] we included a self-contained proof of this fact, which makes it apparent that there is an open subset $U=U_{I}$ of $\mathbb{C}$} which contains $I$, independent of $f$, such that the function $g$ in the conclusion of the theorem can be taken as the restriction of a holomorphic function $U\to{\mathbb{C}}$}. If $I$ is bounded, then the open set $U$ obtained through the proof of Whitney’s theorem as in loc. cit. is also bounded. (See [3, Corollary A.4].) However, if $I={\mathbb{R}}$, then this procedure yields $U={\mathbb{C}}$, so $g$ can be taken as the restriction of an entire function. Independently of Whitney, this fact had also been discovered by Hoischen [14]; for $r=0$, this is Carleman’s 1927 theorem [8] mentioned above, and for $r=1$ this was shown by Kaplan [15] in 1955. (See [21] for the history of such approximation theorems, going back to Weierstrass.) By a theorem of Gauthier and Kienzle [13, Theorem 1.3], in the case $r<\infty$} one can always take $U=I\cup({\mathbb{C}}\setminus{\mathbb{R}})$. Our first main result strengthens this fact by showing that it is possible to choose $U={\mathbb{C}}\setminus\{\alpha,\beta\}$ (even when $r=\infty$):
@@
If $\lambda_{m}(g)<\infty$ for some $m$, then $g$ is said to have finite index, and in this case the smallest such $m$ is called the index of $g$. Note that if $g$ has finite index $m$, then $m\geqslant 1$}. By convention, constant functions $\mathbb{C}$$\to{\mathbb{C}}$ have index 0. The entire functions of index $1$ are exactly the non-constant polynomial functions. The entire functions of index $\leqslant 2$ are also known as the entire functions of finite order, and  $\lambda_{2}(g)$ is called the order of $g$. (See [6, Chapter 2].) By [24, Theorem 1], if $g$ has index $m\geqslant 2$, then $\lambda=\lambda_{m}(g)$ may be computed from the Taylor coefficients $g_{n}:=g^{(n)}(0)/n!$ of $g$ at 0 as follows:
@@
is the basis of a topology on $\mathcal{S}({\mathbb{R}})$, which makes the $\mathbb{R}$}-vector space $\mathcal{S}({\mathbb{R}})$ into a topological $\mathbb{R}$}-vector space (in fact, a Fréchet space), known as the Schwartz space on $\mathbb{R}$}, which plays an important role in the theory of distributions (cf. [25]). It is well-known that the set of rapidly decreasing functions which extend to entire functions is dense in $\mathcal{S}({\mathbb{R}})$; see [25, Theorem 15.5, p. 160]. From Theorem 2 we obtain an effective version of this fact:
@@
Acknowledgements
The author would like to thank Paul Gauthier and Herwig Hauser for various suggestions to improve this paper.
@@ 2. Proof of Theorem 1
Next we prove Corollary 1. Let $I$, $f$, $\varepsilon$, $\rho$ be as in the corollary. The case where $I$ is open is covered by Theorem 1, and the case where $I$ is closed and bounded is a consequence of the Weierstrass Approximation Theorem [19, p. 33]. Suppose $I=[\alpha,\beta)$} where $\infty<\alpha<\beta\leqslant\infty$. (The remaining case where $I=(\alpha,\beta]$ with $\infty\leqslant\alpha<\beta<\infty$ is treated in a similar way.) Then $\operatorname{fr}(I)=\{\beta\}$. We set $I_{1}:=(-\infty,\beta)$ and now use a well-known fact:
@@ 5. Whitney Approximation with Bounds
$\leqslant\big{(}2c(\rho(s)+1)\big{)}^{d(\rho(s)+1)(s/\sqrt{2}-1)}\cdot\lVert f\rVert_{s;\,\rho(s)+1}.$